Find the magnitude of the electric field of a large plate with surface charge density of 300 m at a point 2.00 m away from the surface? 7.
Answer #1

Given surface charge density

\sigma=300\frac{c}{m^2}

Magnitude of electric field is

E=\frac{\sigma}{2\epsilon_o}=\frac{300}{2*8.85*I0^{-12}}V/m

E=1.695*10^{13}V/m